OpenStudy (baby456):

For the equation of a circle x^2+y^2+6y=27, complete the square to put it in the form(x-h)^2+(y-k)^2=r^2, then give center and radius.

OpenStudy (baby456):

@zepdrix

OpenStudy (baby456):

@EmmaMink

OpenStudy (baby456):

@jim_thompson5910

OpenStudy (jdoe0001):

have you covered "perfect square trinomials" yet? or know what a perfect square trinomial is

OpenStudy (baby456):

i forgot it.

OpenStudy (emmamink):

hmm

OpenStudy (emmamink):

It's been a while since I've done this

OpenStudy (jdoe0001):

well, you may want to brush up on them before doing the exercise then check your book for the section on "perfect square trinomial" or sometimes also called "perfect square"

OpenStudy (jdoe0001):

would look like \(\begin{array}{cccccllllll} {\color{brown}{ a}}^2& - &2{\color{brown}{ a}}{\color{blue}{ b}}&+&{\color{blue}{ b}}^2\\ \downarrow && &&\downarrow \\ {\color{brown}{ a}}&& &&{\color{blue}{ b}}\\ &\to &({\color{brown}{ a}} - {\color{blue}{ b}})^2&\leftarrow \end{array}\)

OpenStudy (jdoe0001):

or \(\begin{array}{cccccllllll} {\color{brown}{ a}}^2& + &2{\color{brown}{ a}}{\color{blue}{ b}}&+&{\color{blue}{ b}}^2\\ \downarrow && &&\downarrow \\ {\color{brown}{ a}}&& &&{\color{blue}{ b}}\\ &\to &({\color{brown}{ a}} + {\color{blue}{ b}})^2&\leftarrow \end{array}\qquad % perfect square trinomial, negative middle term \begin{array}{cccccllllll} {\color{brown}{ a}}^2& - &2{\color{brown}{ a}}{\color{blue}{ b}}&+&{\color{blue}{ b}}^2\\ \downarrow && &&\downarrow \\ {\color{brown}{ a}}&& &&{\color{blue}{ b}}\\ &\to &({\color{brown}{ a}} - {\color{blue}{ b}})^2&\leftarrow \end{array}\) for the positive or negative one

jimthompson5910 (jim_thompson5910):

this might help http://www.regentsprep.org/regents/math/algtrig/ate12/completesq.htm

OpenStudy (baby456):

ok